An herbalist has 30 ounces of herb costing $2.00 per ounce. How many ounces of herbs costing $1.00 per ounce should be mixed with the 30 ounces to produce a misture costing $1.60 per ounce?
Answer by nerdybill(7384)

An herbalist has 30 ounces of herb costing $2.00 per ounce. How many ounces of herbs costing $1.00 per ounce should be mixed with the 30 ounces to produce a misture costing $1.60 per ounce?
.
Let x = amount of $1.00 per ounce herb added
.
2(30) + 1(x) = 1.60(30+x)
60 + x = 48 + 1.60x
60 = 48 + .60x
12 = .60x
20 ounces = x (amt added)